A solid cylinder, 0 ≤ r ≤ b, is initially at zero temperature. For times t > 0, internal energy is generated in the solid at a rate of g(t) per unit volume (W/m3), whereas the boundary surface at r = b dissipates heat by convection into a medium at zero temperature. Obtain an expression for the temperature distribution T(r,t) in the cylinder for times t > 0.



1
Expert's answer
2021-12-20T13:58:02-0500

At "t>0" "r=0,T(r,t)=\frac{g(t)}{v}"

At "t>0" "r=b,T(r,t)=-g(t)"


"\therefore\ T(r,t)=\frac{g(t,)}{v}-g(t)"